# Synomaly Noise and Multi-Stage Diffusion: A Novel Approach for Unsupervised Anomaly Detection in Ultrasound Imaging

### v1

@@ Abstract
Ultrasound (US) imaging is widely used in routine clinical practice due to its advantages of being radiation-free, cost-effective, and portable. However, the low reproducibility and quality of US images, combined with the scarcity of expert-level annotation, make the training of fully supervised segmentation models challenging. To address these issues, we propose a novel unsupervised anomaly detection framework based on a diffusion model that incorporates a synthetic anomaly (Synomaly) noise function and a multi-stage diffusion process. Synomaly noise introduces synthetic anomalies into healthy images during training, allowing the model to effectively learn anomaly removal. The multi-stage diffusion process is introduced to progressively denoise images, preserving fine details while improving the quality of anomaly-free reconstructions. The generated high-fidelity counterfactual healthy images can further enhance the interpretability of the segmentation models, as well as provide a reliable baseline for evaluating the extent of anomalies and supporting clinical decision-making. Notably, the unsupervised anomaly detection model is trained purely on healthy images, eliminating the need for anomalous training samples and pixel-level annotations. We validate the proposed approach on carotid US, brain MRI, and liver CT datasets. The experimental results demonstrate that the proposed framework outperforms existing state-of-the-art unsupervised anomaly detection methods, achieving performance comparable to fully supervised segmentation models in the US dataset. Additionally, ablation studies underline the importance of hyperparameter selection for Synomaly noise and the effectiveness of the multi-stage diffusion process in enhancing model performance. Code: https://github.com/yuan-12138/Synomaly.
@@
Ultrasound Image Analysis, Unsupervised Anomaly Detection, Diffusion Model, Medical Image Analysis, Carotid imaging
@@ 1 Introduction
Ultrasound (US) plays a crucial role in daily medical practice. It is commonly used for the screening and early assessment of internal organs and vasculature. Compared to computed tomography (CT) and magnetic resonance imaging (MRI), US is more frequently utilized for routine examinations. CT, although efficient in providing high-quality images, involves radiation exposure, which makes it less suitable for regular check-ups. MRI, while providing detailed soft tissue contrast, is often cost- and time-intensive for routine use. Consequently, physicians typically reserve CT and MRI for cases where a patient has already exhibited specific symptoms or conditions. As a result, CT and MRI are more commonly associated with the imaging of diseased cases. In contrast, US is radiation-free, cost-effective, and portable, making it an ideal modality for regular monitoring. Therefore, the US images collected are often from non-diseased individuals, making it as an ideal modality to learn the physiology of healthy tissues.
However, the inherent challenges of low reproducibility and poor image quality in US imaging make it difficult to gather large amounts of diseased data for training of deep neural networks. This challenge is particularly evident in conditions like stroke, a leading cause of mortality worldwide, with ischaemic stroke accounting for about 65.3% of all cases [16]. This type of stroke is caused by the blockage of a blood vessel supplying the brain, and carotid atherosclerosis is a common underlying cause [22]. US is frequently used for the detection and monitoring of carotid plaques. One important indicator to assess the risk of stroke for atherosclerosis patients is total plaque volume [1]. One pre-requisite for such measurement is an accurate segmentation of the carotid plaque in US images. However, unlike healthy anatomy, plaques vary significantly in shape and location, making their detection more challenging. This variability means that high-quality datasets are crucial for training segmentation models to reliably identify plaques. Without such datasets, models may struggle to generalize and correctly segment plaques in diverse cases. Because of ethical and privacy issues, it is not an easy practice to gather large amount of diseased US data. To tackle such challenge, unsupervised anomaly detection strategy appears to be an effective solution.
@@
Similar to how humans learn about diseased tissues, unsupervised anomaly detection method tends to distinguish the anomaly tissue from the normal anatomy by learning from healthy structures [17]. Generative models are commonly utilized to generate synthetic counterfactual healthy samples based on diseased images. Since the model is purely trained on healthy data, the need for pixel-level labels is eliminated. Considering the high noise ratio and low contrast of US compared to CT and MRI, the labelling process for US sometimes introduces unavoidable ambiguities, thus complicating the fully-supervised training process [15]. However, such drawback does not concern the unsupervised anomaly detection method. Furthermore, it also provides the possibility to detect unseen or rare anomaly cases, which are not well represented in the dataset for fully-supervised models. Nonetheless, unsupervised anomaly detection is less common in US due to the presence of speckle noise, shadows, and acoustic artifacts, which make it challenging to differentiate anomalies from healthy tissues. Such complexity of US images necessitates the development of more advanced generative models.
@@
Diffusion models have proven effective in modeling complex data distributions, resulting in higher image sample quality compared to generative adversarial networks (GANs) and variational autoencoders (VAEs) [12, 10]. These advantages are particularly valuable for unsupervised anomaly segmentation, especially in cases of small datasets, leading to growing interest in applying diffusion models to this application domain [27]. Most recent works have also reported initial attempts of applying diffusion models for unsupervised anomaly detection in US [34, 32], specifically exploring their effectiveness for detecting anomalies in fetal brain US. The performance of diffusion model is closely linked to the choice of corrupting noise, especially for medical image anomaly detection [26]. Superior performance over naive pixel-wise Gaussian noise has been reported when applying Simplex noise [42] and Coarse noise [25] for anomaly detection of brain MRI.
@@
In order to develop a robust unsupervised anomaly detection model for US images, in this work, we propose Synomaly noise, which stands for synthetic anomaly noise based on statistical disease morphology prior. Different from the aforementioned noise functions, Synomaly noise is specifically designed for unsupervised medical anomaly detection by corrupting healthy images with synthetic anomalies during training. This design of noise can result in a better erasure of anomalies in the images. To further suppress the loss of detailed information during noising process, a multi-stage diffusion process is proposed for inference. The detailed contributions are summarized as follows:
@@
This is the first work to apply a diffusion model for unsupervised anomaly detection in carotid US images. We demonstrate the effectiveness of the proposed Synomaly noise and multi-stage diffusion process on a US carotid plaque dataset. To further validate the performance of the proposed approach, we conducted experiments on public brain MRI and liver CT datasets, achieving superior performance over state-of-the-art methods in detecting brain tumors and liver diseases, respectively.
The results of the three most common medical image modalities (US, CT, and MRI) demonstrate the potential of the proposed method in wide application. To facilitate the reproducibility of this work, the code will be released on this website upon acceptance.
@@ 2.1 Unsupervised Anomaly Detection Using Autoencoders
Autoencoders (AEs), since their introduction, have been widely used for unsupervised representation learning. The principles behind AE-based unsupervised anomaly detection/segmentation is to learn a latent space distribution that represents healthy anatomies, identifying abnormalities as outliers. Consequently, the reconstructed images preserve the healthy regions while altering the anomalous parts, making it easier to detect and segment the abnormalities. Different variants of AEs have been applied to different anomaly detection scenarios. [40] implemented a conditional VAE to perform anomaly segmentation in brain MRI, while [35] utilized a Bayesian convolutional AE to locate the abnormalities in CT. By combining VAE with a self-supervised image representation learning network, [29] proposed SSL-AnoVAE to tackle the anomaly staging and segmentation of retinal images. Deviated from the reconstruction training approaches, [44] adopted an inpainting strategy to train a context-encoding VAE for anomaly detection in brain MRI. In order to increase the ability of AEs in modeling complex high-dimensional data distributions, adversarial strategies have been introduced [21]. [5] integrated GAN with VAE and proposed AnoVAEGAN to perform anomaly detection for brain MRI.
Apart from reconstruction-based approaches, restoration-based methods offer an alternative. Instead of directly passing input images through AE network to generate counterfactual reconstructions, restoration methods focus on finding the closest representation of the abnormal image in the learned latent space of normal images. These methods project the abnormal image onto the normal manifold and transform it into its closest healthy version. Anomalies are then detected by comparing the original input with the restored image. [39] proposed AnoGAN, where a GAN is trained on healthy images to model the latent manifold for normal images. During inference, a random latent point is sampled and iteratively moved towards the closest counterpart on the normal manifold using a gradient decent method. In an extended work, [38] implemented an encoder network to replace the time-consuming iterative restoration process. [11] trained a VAE to model the normative data distribution. The restoration process of anomaly samples is performed through an iterative gradient ascent method by Maximum-A-Posteriori estimation.
@@ 2.2 Unsupervised Anomaly Detection Using Denoising Models
Recent advances in diffusion models have demonstrated superior generative performance across various medical imaging modalities [13, 43, 14, 24]. This enhanced performance has also proven beneficial for unsupervised anomaly detection in medical imaging, enabling more accurate identification of abnormalities [27, 28]. [42] proposed AnoDDPM, where a Denoising Diffusion Probabilistic Model (DDPM) is utilized to reverse the diffusion process and generate counterfactual images. A partial diffusion process is introduced to better conserve the fine details of the input images and further increase the sample time efficiency during inference. They also replaced the common Gaussian noise with Simplex noise for image corruption and achieved better results when testing on brain MRI images. To avoid the accumulated errors during iterative denoising process and increase the sample efficiency of the diffusion process, [36] implemented latent diffusion model for anomaly detection in brain MRI and CT. The noising and denoising process is then executed in the latent space rather than directly on the images, which can significantly lower the dimensionality and improve efficiency. [41] utilized a Denoising Diffusion Implicit Model for medical anomaly detection, which enables a much faster sampling than DDPMs. The diffusion model is trained on both healthy and diseased images, and during inference, the generation of healthy counterparts for diseased images is guided by a pretrained classifier. Instead of providing direct guidance during the denoising process, [8] introduced THOR to provide implicit guidances through intermediate anomaly maps. This approach allows the model to focus on potential anomaly regions while preserving more details in the healthy areas, thereby reducing the false positive rate in anomaly detection. Similarly, [19] combined DDPM with ACAT [18], a saliency map that can approximately capture the pathological regions in the input image. At each denoising step, the regions identified by the ACAT saliency map are processed separately enabling targeted modification of potential diseased areas while preserving healthy regions. As a result, they demonstrated enhanced erasure performance of the anomalous regions in experiments on brain MRI and CT datasets. Considering the high computational burden of the diffusion process for high-resolution images like MRI, [6] proposed a patch-based diffusion model, which reconstructs input images patchwise, thereby reducing network complexity.
@@ 3 Method
The performance of diffusion models in unsupervised anomaly detection is closely related to the design of the noising function. An effective noise function should efficiently corrupt the original image such that, after the noising process, the anomalous regions become indistinguishable from the normal ones. This allows the model to replace anomalies with normal tissues during the denoising process. However, the classic noise functions face challenges with larger or particularly distinct anomalies, such as very bright or dark regions. In such cases, many noise steps may be required to sufficiently blur the anomaly and make it indistinguishable from the surrounding healthy tissue. This extensive noise addition can lead to significant loss of structural image information, resulting in blurred and inaccurate image reconstruction. This introduces a precision-recall trade-off: Using fewer noise steps preserves structural integrity and achieves high reconstruction accuracy for healthy tissues, but it increases the risk of failing to detect distinct anomalies, leading to a higher rate of false negatives - characterized by high precision and low recall. On the other hand, using more noise steps enhances the likelihood of detecting anomalies, but it also raises the possibility of false positives, where healthy tissues are incorrectly identified as anomalies - characterized by low precision and high recall.
To tackle these challenges, we propose Synomaly noise. The core idea is to corrupt healthy images by introducing synthetic anomalous regions during training, simulating the appearance of noised anomalous regions, as shown in Fig. 1(a). This allows the model to refine its denoising strategy specifically for reconstructing normal tissue from noised regions that resemble pathological anomalies. During inference, standard Gaussian noise is applied to the images, causing the abnormal regions to resemble the noised artificial anomalies encountered during training. This similarity enables the model to efficiently generate counterfactual healthy tissues and effectively replace anomalies with normal tissue during the denoising process.
@@ 3.2 Synomaly Noise
Synomaly noise stands for synthetic anomaly noise. Instead of merely adding noise until anomalies blend into healthy tissues, Synomaly strategically integrates synthetic anomalies into the noise model during training. The core concept is to train the diffusion model to remove synthetic anomalies in healthy images that mimic real-world scenarios, thereby improving the model’s ability to detect and remove actual anomalies during inference. The key advantage of Synomaly in comparison to the other noise functions is its ability to utilize disease-specific information (morphology or appearance of the disease) during model training, which greatly enhances anomaly detection performance based on our experiments.
Fig. 2 illustrates the process of Synomaly noise generation. In the first step, a pure Gaussian noise is created, matching the dimensions of the input image. This Gaussian noise will also serve as the Gaussian background noise of Synomaly noise. Next, the shapes of the synthetic anomalies are generated by applying Gaussian blurring to the initial noise, followed by thresholding. Here, the variance of Gaussian blurring () as well as the threshold value () are predefined as model training parameters and control the size and number of synthetic anomalies. Specifically, increasing the blurring variance () results in larger fragments of anomalies, while higher threshold values () reduce the number and size of anomalies by excluding lower Gaussian values. To further refine the simulation of real anomalies, an optional mask can be applied to constrain the anomaly shapes to relevant regions.
@@ 3.3 Multi-stage Diffusion Process at Inference
Most of the existing works on diffusion models for anomaly detection typically suggest performing the diffusion process only once during inference, which is termed as “single-stage” approach. However, this approach is usually insufficient when dealing with anomalies that are unique in shape or intensity. Often, a single-stage cycle cannot fully remove anomalies, especially when many noise steps are required to achieve good anomaly detection. While increasing the number of noise steps can improve anomaly removal, it can also degrade structural information and eventually lead to blurred and inaccurately reconstructed images.
Therefore, we propose a “multi-stage diffusion process”. Instead of applying many noise steps in a single inference stage, the multi-stage approach repeatedly applies a few noise steps followed by denoising across multiple inference stages. As depicted in Fig. 3(a), starting with the original input image, the process begins by adding noise for a predefined number of noise steps, followed by denoising. Then, the denoised image is fused with the original image to have the image with reduced abnormality for the next stage. This fusion process is detailed in Fig. 3(b), where the computed anomaly mask in the current stage is used to provide guidance. This process is repeated over multiple stages until the generated anomaly map converges. To prevent excessive iterations, the maximum number of iterations is set to five.
This approach offers several benefits. By using fewer noise steps at each stage, the structural integrity of the image is better preserved, leading to more accurate reconstructions. Additionally, performing multiple inference stages with sequential denoising allows even distinct anomalies to be progressively removed and detected more effectively. Since different noise steps model different image features [23] and the goal of medical anomaly detection is to focus on finer details without excessively corrupting the entire image, fewer noise steps are particularly relevant for detecting subtle anomalous regions.

### v2

# Synomaly Noise and Multi-Stage Diffusion: A Novel Approach for Unsupervised Anomaly Detection in Medical Images

@@ Abstract
Anomaly detection in medical imaging plays a crucial role in identifying pathological regions across various imaging modalities, such as brain MRI, liver CT, and carotid ultrasound (US). However, training fully supervised segmentation models is often hindered by the scarcity of expert annotations and the complexity of diverse anatomical structures. To address these issues, we propose a novel unsupervised anomaly detection framework based on a diffusion model that incorporates a synthetic anomaly (Synomaly) noise function and a multi-stage diffusion process. Synomaly noise introduces synthetic anomalies into healthy images during training, allowing the model to effectively learn anomaly removal. The multi-stage diffusion process is introduced to progressively denoise images, preserving fine details while improving the quality of anomaly-free reconstructions. The generated high-fidelity counterfactual healthy images can further enhance the interpretability of the segmentation models, as well as provide a reliable baseline for evaluating the extent of anomalies and supporting clinical decision-making. Notably, the unsupervised anomaly detection model is trained purely on healthy images, eliminating the need for anomalous training samples and pixel-level annotations. We validate the proposed approach on brain MRI, liver CT datasets, and carotid US. The experimental results demonstrate that the proposed framework outperforms existing state-of-the-art unsupervised anomaly detection methods, achieving performance comparable to fully supervised segmentation models in the US dataset. Ablation studies further highlight the contributions of Synomaly noise and the multi-stage diffusion process in improving anomaly segmentation. These findings underscore the potential of our approach as a robust and annotation-efficient alternative for medical anomaly detection. Code: https://github.com/yuan-12138/Synomaly.
@@
[…]
@@ 1 Introduction
Anomaly detection in medical imaging is a common task in medical image analysis, where pathological regions are identified for further evaluation and classification. For example, detecting suspicious tumor regions in the brain and liver from magnetic resonance imaging (MRI) or computed tomography (CT) scans, or identifying carotid plaques in ultrasound (US) images, is essential for assessing disease severity and guiding clinical decisions. Traditionally, these tasks have been performed manually by radiologists and sonographers—a process that is time-consuming, labor-intensive, and prone to inter-operator variability.
With recent advances in deep learning, automated image analysis has emerged as a powerful tool to assist medical professionals in detecting and evaluating pathologies more efficiently [30, 51, 29, 11, 27, 32]. However, the effectiveness and robustness of deep learning models depend heavily on training data quality. In the natural image domain, data collection is relatively straightforward, with images readily available online or captured using consumer-grade cameras. Annotations can be performed by large groups of non-experts, enabling rapid labeling of extensive datasets at a low cost. In contrast, medical imaging data is restricted by privacy regulations and ethical considerations, requiring institutional approval and de-identification [31]. Annotation is costly and time-consuming, as it demands specialized expertise from physicians [12]. Additionally, inter-operator variability in labeling can introduce bias, further limiting model generalization. To tackle this challenge, unsupervised anomaly detection approach appears to be an effective solution.
@@
Similar to how humans learn about diseased tissues, unsupervised anomaly detection method tends to distinguish the anomaly tissue from the normal anatomy by learning from healthy structures [21]. Generative models are commonly utilized to synthesize counterfactual healthy samples from diseased images, enabling anomaly detection without requiring pixel-level annotations [15]. This eliminates the need for extensive manual labeling, which can be challenging due to variations in image quality, contrast, and noise across different imaging modalities. For example, the high noise ratio and low contrast of US make the labelling process for US sometimes introduces unavoidable ambiguities, thus complicating the fully-supervised training process [20]. However, such drawback does not concern the unsupervised anomaly detection method. Furthermore, it also provides the possibility to detect unseen or rare anomaly cases, which are not well represented in the dataset for fully-supervised models. Despite its advantages, the effectiveness of unsupervised anomaly detection largely depends on the generative model’s ability to preserve healthy structures and accurately reconstruct pathological regions as their healthy counterparts. This underscores the need for further exploration of advanced generative models to enhance performance.
@@
Diffusion models have proven effective in modeling complex data distributions, resulting in higher image sample quality compared to generative adversarial networks (GANs) and variational autoencoders (VAEs) [17, 14, 43]. These advantages are particularly valuable for unsupervised anomaly segmentation, especially in cases of small datasets, leading to growing interest in applying diffusion models to this application domain [35]. Recent works have also reported several attempts of applying diffusion models for unsupervised anomaly detection in fetal brain US [47, 45], brain MRI [40], and brain CT [23]. The performance of diffusion model is closely linked to the choice of corrupting noise, especially for medical image anomaly detection [34]. Superior performance over naive pixel-wise Gaussian noise has been reported when applying Simplex noise [61] and Coarse noise [33] for anomaly detection of brain MRI.
@@
In order to develop a robust unsupervised anomaly detection model for medical images, in this work, we propose Synomaly noise, which stands for synthetic anomaly noise based on statistical disease morphology prior. Different from the aforementioned noise functions, Synomaly noise is specifically designed for unsupervised medical anomaly detection by corrupting healthy images with synthetic anomalies during training. This design of noise can result in a better erasure of anomalies in the images. To further suppress the loss of detailed information during noising process, a multi-stage diffusion process is proposed for inference. The detailed contributions are summarized as follows:
@@
We conduct extensive validation across multiple medical imaging modalities, including brain MRI, liver CT, and carotid ultrasound (US) datasets. To the best of our knowledge, this is the first work to apply a diffusion model for unsupervised anomaly detection in carotid US images. Our approach achieves superior performance compared to state-of-the-art methods.
The results of the three most common medical image modalities (US, CT, and MRI) demonstrate the potential of the proposed method in wide application. To facilitate the reproducibility of this work, the code will be released on this website upon acceptance.111Codes: https://github.com/yuan-12138/Synomaly
@@ 2.1 Unsupervised Anomaly Detection Using Autoencoders
Autoencoders (AEs) have been widely used for unsupervised anomaly detection by learning latent representations of healthy anatomy and identifying deviations as anomalies. The reconstructed outputs typically preserve normal regions while modifying anomalous ones, aiding detection and segmentation. Various AE variants have been applied across modalities [1, 41, 55]. For example, [59] used a conditional VAE for brain MRI segmentation, and [48] applied a Bayesian convolutional AE to CT. [39] proposed SSL-AnoVAE by combining VAE with self-supervised learning for retinal images. Instead of standard reconstruction, [64] adopted a context-encoding VAE trained via inpainting. To improve AE capacity for modeling complex distributions, adversarial methods have been introduced. [5] combined VAE with GAN and proposed AnoVAEGAN for brain MRI anomaly detection.
Apart from reconstruction-based approaches, restoration-based methods offer an alternative. Instead of directly passing input images through AE network to generate counterfactual reconstructions, restoration methods focus on finding the closest representation of the abnormal image in the learned latent space of normal images. By projecting the abnormal image onto the normal manifold, the model transforms it into its most similar healthy counterpart [25]. Anomalies are then detected by comparing the original input with the restored image. [53] proposed AnoGAN, where a GAN is trained on healthy images to model the latent manifold for normal images. During inference, a random latent point is sampled and iteratively moved towards the closest counterpart on the normal manifold using a gradient decent method. In an extended work, [52] implemented an encoder network to replace the time-consuming iterative restoration process. [16] trained a VAE to model the normative data distribution. The restoration process of anomaly samples is performed through an iterative gradient ascent method by Maximum-A-Posteriori estimation.
2.2 Unsupervised Anomaly Detection Using Synthetic Anomalies
Another mainstream direction in unsupervised anomaly detection is the synthetic anomaly approach, where the goal is to generate anomalous-looking images from normal ones to train a discriminative model. For example, CutPaste [37] creates pseudo-anomalies by cutting and pasting patches within a normal image to simulate realistic-looking defects. These synthesized images are then used to train a model to directly predict anomaly maps or scores. To improve realism, Natural Synthetic Anomalies (NSA) [54] employ Poisson image editing to seamlessly blend scaled patches from different images, enhancing the variability and natural appearance of the simulated anomalies. Similarly, in the medical domain, [58] applied this paradigm to chest X-ray and US images using Poisson image interpolation, generating anomalies that mimic real pathological structures. However, these methods do not produce counterfactual reconstructions and are dependent on the fidelity and realism of the generated pseudo-anomalies [57]. If the true anomalies have highly out-of-distribution appearances, these approaches may struggle to generalize. Fundamentally, such methods attempt to model what anomalies might look like, rather than learning what healthy tissue should look like, as is done in reconstruction-based methods.
@@ 2.3 Unsupervised Anomaly Detection Using Denoising Models
Recent advances in diffusion models have shown strong generative performance across medical imaging modalities [18, 62, 19, 28], which has also benefited unsupervised anomaly detection by enabling more accurate identification of abnormalities [35, 38]. [61] introduced AnoDDPM, a DDPM-based method for generating counterfactual images using a partial diffusion process to preserve fine image details and improve inference speed. They also adopted Simplex noise over Gaussian for better results on brain MRI. To avoid accumulated errors and improve sampling efficiency, [49] proposed a latent diffusion model, operating in compressed latent space rather than image space. [60] used a Denoising Diffusion Implicit Model, enabling faster inference. Their model was trained on both healthy and diseased images, with generation guided by a pretrained classifier. [9] proposed THOR, which incorporates anomaly maps during inference to focus on suspicious regions while preserving healthy structures. [23] combined DDPM with ACAT [22], using saliency maps to isolate and process potential anomalies separately, improving anomaly erasure in brain MRI and CT. To address the high computational cost for high-resolution images like MRI, [6] proposed a patch-based diffusion model that reconstructs images patchwise, reducing complexity.
@@ 3 Method
The performance of diffusion models in unsupervised anomaly detection is closely tied to the design of the noising function. An effective noise function should corrupt the image such that, after noising, anomalous regions become indistinguishable from normal tissue—allowing the model to replace them with healthy structures during denoising. However, classic noise functions struggle with large or highly distinct anomalies, e.g., very bright or dark regions, which often require many noise steps to obscure. This heavy corruption can degrade structural details and lead to blurred, inaccurate reconstructions. This introduces a precision-recall trade-off: Using fewer noise steps preserves structural integrity and achieves high reconstruction accuracy for healthy tissues, but it increases the risk of failing to detect distinct anomalies, leading to a higher rate of false negatives - characterized by high precision and low recall. On the other hand, using more noise steps enhances the likelihood of detecting anomalies, but it also raises the possibility of false positives, where healthy tissues are incorrectly identified as anomalies - characterized by low precision and high recall.
To tackle these challenges, we propose Synomaly noise. The core idea is to corrupt healthy images by introducing synthetic anomalous regions during training, simulating the appearance of noised anomalous regions, as shown in Fig. 1(a). This enables the model to refine its denoising strategy for reconstructing normal tissue from corrupted regions resembling real corrupted anomalies. During inference, standard Gaussian noise is applied, making abnormal areas resemble the synthetic anomalies seen during training. This similarity allows the model to efficiently generate counterfactual healthy tissue and replace anomalies during denoising. Importantly, the goal of Synomaly noise is not to produce anatomically realistic lesions, but to create synthetic anomalies that mimic the visual appearance of real anomalies after noise corruption.
@@ 3.2 Synomaly Noise
Synomaly noise stands for synthetic anomaly noise. Instead of merely adding noise until anomalies blend into healthy tissues, Synomaly strategically integrates synthetic anomalies into the noise model during training. The core idea is to train the diffusion model to remove synthetic anomalies embedded in healthy images that resemble real pathological patterns, thereby enhancing its ability to detect and remove true anomalies at inference.
Fig. 2 illustrates the process of Synomaly noise generation. In the first step, a pure Gaussian noise is created, matching the dimensions of the input image. This Gaussian noise will also serve as the Gaussian background noise of Synomaly noise. Next, the shapes of the synthetic anomalies are generated by applying Gaussian blurring to the initial noise, followed by thresholding. Here, the variance of Gaussian blurring () as well as the threshold value () are predefined as model training parameters and control the size and number of synthetic anomalies. Specifically, increasing the blurring variance () results in larger fragments of anomalies, while higher threshold values () reduce the number and size of anomalies by excluding lower Gaussian values. To further refine the simulation of real anomalies, an optional mask can be applied to ensure that synthetic anomalies are placed within relevant anatomical regions rather than in background areas. This mask is particularly useful for imaging modalities where the foreground and background are not clearly separable or where anomalies are expected to occur within specific structures.
@@ 3.3 Multi-stage Diffusion Process at Inference
Most of the existing works on diffusion models for anomaly detection typically suggest performing the diffusion process only once during inference, which is termed as “single-stage” approach. However, this is often inadequate for anomalies with distinct appearance. A single pass may not fully remove such anomalies, especially when many noise steps are needed. While increasing noise steps can improve anomaly removal, it also degrades structural detail, resulting in blurred and inaccurate reconstructions.
Therefore, we propose a “multi-stage diffusion process”. Instead of applying many noise steps in a single inference stage, the multi-stage approach repeatedly applies a few noise steps followed by denoising across multiple inference stages. As depicted in Fig. 3(a), starting with the original input image , the process begins by adding noise for a predefined number of noise steps, followed by denoising. Based on the denoised image , the anomaly mask is generated (Fig. 3(b)) by thresholding the difference between blurred versions of the denoised and the original image. Using this mask, masked fusion (Fig. 3(c)) combines anomalous regions from with the background from the original image , producing a new input for the next stage. This process iterates until the sum of detected anomaly pixels stabilizes, meaning that the relative change between consecutive anomaly masks falls below a threshold of 0.01. As a result, the number of diffusion stages at inference is not a manually set hyperparameter but is automatically determined based on the convergence of the anomaly mask. To prevent excessive computations, a maximum of five iterations is imposed, as further iterations typically yield negligible improvements.
This approach offers several benefits. By decomposing a large noise step into multiple smaller steps across stages, the structural integrity of the image is better preserved, leading to more accurate reconstructions. Additionally, performing multiple inference stages with sequential denoising allows even distinct anomalies to be progressively removed and detected more effectively. Since different noise steps model different image features [26] and the goal of medical anomaly detection is to focus on finer details without excessively corrupting the entire image, fewer noise steps are particularly relevant for preserving details from non-anomalous regions.